What is the effective resistance between BC if the value of each resistor, R, is 120 Ω?BRD E C A
Answer #1

Here, for the resistance between B and C

for R and R in series

RCD = 2R

for 2R and R in parallel

1/Req = 1/(2R) + 1/R

Req = 2R/3

NOw, R and Req in series

R3 = 2R/3 + R = 5R/3

for R3 , 2R , R in parallel

1/Req = 1/(5R/3) + 1/(2R) + 1/R

1/Req = 1/(5 * 120/5) + 1/(2 * 120) + 1/120

Req = 48 Ohm

the effective resistance is 48 Ohm